# Compact Speech Translation Models via Discrete Speech Units Pretraining

### v1

@@ 5.1 Tokenization and its Transfer
When BPE is not applied on the DSU, the DSU-to-Trl models have a BLEU of $6.48$$\pm$ $0.26$ across the 6 vocabulary settings. Despite being smaller (<$30$M), these models are better than the Scratch model of $5.99$ BLEU.
When BPE is applied, the DSU sequence length becomes a crucial factor. On one hand, it delivers the best model ("$8$K": $7.02$ BLEU). On the other hand, the result deteriorates a lot in the other 5 settings, i.e., a BLEU of $5.12$$\pm$ $0.83$. More importantly, BPE’s correlation with the model performance is not straightforward, neither is the length ratio. For an example, the "$32$K" model (DSU length of $118$) is about $2$ BLEU points behind to the "$8$K" model (DSU length of $150$).
DSU-Adapter is more tokenization robust
@@
When BPE is applied on the DSU, DSU-Adapter has a higher BLEU of $8.86$$\pm$ $0.19$, in contrast to $8.17$$\pm$ $0.32$. This behaviour is opposed to the DSU-to-Trl model that has a lower BLEU of $5.54$$\pm$ $1.07$ (with also larger variance), in contrast to $6.48$$\pm$ $0.26$ when BPE is not used. The better scores with smaller variance show that DSU-Adapter is more tokenization robust. We see this as a benefit of introducing the DSU (SSL knowledge) via PT rather than as model inputs. Besides, using separate vocabularies in PT is preferred since it performs slightly better, and the DSU, which are not needed in the ST output, would not occupy the target vocabulary.
Ablation
@@
Having similar size, DSU-Adapter is better than EncDec-Init and Enc-Init. The BLEU over their 12 settings is $8.51$$\pm$ $0.44$, $8.22$$\pm$ $0.51$, and $7.99$$\pm$ $0.61$ respectively. Encoder-initialization seems more crucial than decoder-initialization, as reflected by the fact that the best DSU-Adapter comes from a combination with the weakest DSU-to-Trl model of $4.43$ BLEU.
5.2 CTC regularisation helps
@@
Table 3: Effect of CTC on Fbk-to-DSU PT and/or ST FT to the DSU-Adapter method.
Table 3555Owing to the large number of configurations, we can experiment on 4 vocabularies only and then report their average. presents an analysis of using CTC in the Fbk-to-DSU PT and/or in the ST FT using the DSU-Adapter method. We find that CTC helps on either stage, but the gain is 0.27 BLEU more in finetuning. Using them jointly still helps, but the gain becomes marginal, i.e, 0.03 BLEU.
6 Conclusion
We propose to distill knowledge from the large SSL model via pretraining on its DSU to make ST models more compact for on-device deployment. Our compact model, i.e., the DSU-Adapter, shows stronger result over finetuning the corresponding (larger) HuBERT in multilingual ST task. In contrast to using DSU as model input, our method requires less components in inference and shows stronger robustness over the tokenization. Finally, our method is suitable for the low-resource setting since it requires no transcripts.
7 Limitations
In this work, we propose to distill knowledge from the informative SSL representations via DSU pretraining to make ST models more compact. In addition, we also discuss advantages of our method over using DSU as model inputs. In spite of its noticeable benefits, there are several factors that are not thoroughly explored:
@@ 7 Limitations
We did not inspect the clustering size (fixed to $1,000$) and the number of training instances (only fixed to $98,000$) used in training the K-Means clustering model.
@@
Self-Supervised Models
We did not experiment using other layers (only fixed to the 6th layer) from the HuBERT model nor from stronger SSL models to extract the DSU. These two factors potentially increase the DSU quality and potentially provide better improvement for our method while maintaining the same model size.
@@ Acknowledgements
Appendix A Appendix
A.1 Data preprocessing
log Mel Filterbank
The 80-D log Mel Filterbank is computed every 10ms with a 25ms window, followed by feature normalisation using statistics, i.e., mean and variance, computed over each channel.
BPE
We use the BPE implementation from sentencepiece Kudo and Richardson (2018) and obtain vocabulary of size 8K on the English target, 16K on the transcripts and 32K on the DSU, unless otherwise specified.
HuBERT
K-Means Clustering
We use MiniBatchKMeans from sklearn with a clustering size of $1,000$ and default settings in other hyper-parameters. When training the clustering model, we divide the 21 language pairs into three groups: 1) {ar, cy, et, id, ja, lv, mn, sl, sv, ta, tr}, 2) {nl, pt, ru, zh} and 3) {ca, de, es, fa, fr, it} for sampling. For each language pair in group 1), we sample $1$K instances, whereas it becomes $3$K in group 2) and $12.5$K in group 3), resulting in a total of $98$K instances.
A.2 Model Configuration
We use "pre" layer-normalization Nguyen and Salazar (2019) in DSU-to-Trl.
A.3 Training and Inference
In training,
  • we skip training data that are more than $3$K frames for Filterbank, $320$K frames for HuBERT input or $1,024$ tokens.
  • we set $\lambda_{\alpha}$ and $\lambda_{\beta}$ in CTC to $0.3$ and the smoothing parameter in LS to $0.1$
All models are optimized by Adam Kingma and Ba (2015) with inverse square root scheduler. For simplicity, we provide major training parameters for each model in the format of {effective batch size, warm-up steps, peak learning rate, maximum training steps}:
Fbk-to-Token:
{32K frames, 25K, 2e-3, 60K}
DSU-to-Trl:
{80K tokens, 10K, 5e-4, 50K}
Hu-Transformer:
{4M frames, 4K, 1e-4, 300K}
In inference, we average the last 5 checkpoints and use beam size of 5 in generation. All experiments are run on Nvidia A100 GPUs. It takes about 1 day for 2 A100 (40GB) GPUs to complete an experiment that uses Filterbank as inputs.
A.4 Results and Analysis
In addition to BLEU (see Table 8 for the result of each language), we also provide the result of Table 1 using chrF777nrefs:1—case:mixed—eff:yes—nc:6—nw:0—space:no—version:2.3.1 and COMET. In order to compute the COMET score, we feed the gold-reference transcripts as the source sentence.
Table 4 and Table 6 show the average result in chrF and COMET respectively over the four groups defined in section 5. Their results follow the findings in BLEU: the DSU-Adapter is better than Hu-Transformer and on a par with ASR pretraining. Further details, such as training data size (in terms of the number of utterances rather than hours888”zh” has 10 hours of audio data, whereas ”nl” has only 7 hours. Therefore, ”zh” is classified into ”Mid” resource group even though it has slightly less number of utterances than ”nl”.) and per language performance, are shown on Table 5 for chrF and Table 7 for COMET.
CTC Regularisation
On Table 3, we average the results over this four configurations: 1) "No-BPE 1K-8K", 2) "BPE 8K", 3) "BPE 32K" and 4) "BPE 32K-8K".
AST model Transcripts Model size High Mid Low All
Scratch No 52M $43.58$ $27.22$ $14.62$ $23.10$
ASR Pretraining Yes 52M $51.90$ $32.82$ $16.38$ $27.10$
Hu-Transformer No 113M $49.88$ $31.94$ $16.96$ $26.80$
DSU-Adapter No 48M $52.08$ $33.94$ $16.46$ $27.40$
Table 4: Results of Table 1 in chrF score by resource group. Similar to the findings in BLEU, DSU-Adapter is far better than Hu-Transformer, in particular for "High" and "Mid", and on a par with ASR Pretraining. The best result in each group is denoted by ’_’.
Languages #Training data Resources Scratch ASR-PT Hu-Transformer DSU-Adapter
ca 95,854 High $42.5$ $50.9$ $49.8$ $51.5$
de 127,824 $35.4$ $45.1$ $43.6$ $45.4$
es 79,013 $48.1$ $56$ $52.8$ $55.7$
fr 207,372 $48.3$ $55.6$ $53.3$ $55.7$
fa 53,949 Mid $15.6$ $16.9$ $19.9$ $16.9$
it 31,698 $40.6$ $50.2$ $46.4$ $50.3$
pt 9,158 $25$ $31.3$ $33.8$ $32.3$
ru 12,112 $35.6$ $42.6$ $40$ $47.7$
zh 7,085 $19.3$ $23.1$ $19.6$ $22.5$
ar 2,283 Low $12.5$ $14.4$ $13.8$ $13.2$
cy 1,241 $14.5$ $16.3$ $15.8$ $15.4$
et 1,782 $17.9$ $21.3$ $19$ $19.9$
id 1,243 $13.8$ $14.2$ $14.9$ $14.5$
ja 1,119 $14$ $14$ $13.7$ $14.4$
lv 2,337 $13.9$ $14.5$ $15.9$ $15.4$
mn 2,067 $14.5$ $15.4$ $14.3$ $14.8$
nl 7,108 $17.8$ $24.3$ $30.8$ $27$
sl 1,843 $13.9$ $15.4$ $15.3$ $15.4$
sv 2,160 $13.2$ $13.8$ $17.3$ $14.5$
ta 1,358 $12$ $11.9$ $12$ $12.3$
tr 3,966 $17.4$ $21.1$ $20.7$ $20.7$
Table 5: Results of Table 1 in chrF score for each language pair. The best result in each group is denoted by ’_’.
AST model Transcripts Model size High Mid Low All
Scratch No 52M $0.605$ $0.498$ $0.433$ $0.481$
ASR Pretraining Yes 52M $0.680$ $0.537$ $0.443$ $0.511$
Hu-Transformer No 113M $0.650$ $0.522$ $0.439$ $0.499$
DSU-Adapter No 48M $0.681$ $0.548$ $0.442$ $0.513$
Table 6: Results of Table 1 in COMET score by resource group. Similar to the findings in BLEU, DSU-Adapter is far better than Hu-Transformer, in particular for "High" and "Mid", and on a par with ASR Pretraining. The best result in each group is denoted by ’_’.
Languages #Training data Resources Scratch ASR-PT Hu-Transformer DSU-Adapter
ca 95,854 High $0.607$ $0.681$ $0.654$ $0.684$
de 127,824 $0.535$ $0.608$ $0.588$ $0.613$
es 79,013 $0.620$ $0.696$ $0.654$ $0.692$
fr 207,372 $0.660$ $0.736$ $0.704$ $0.735$
fa 53,949 Mid $0.479$ $0.494$ $0.501$ $0.493$
it 31,698 $0.557$ $0.640$ $0.594$ $0.641$
pt 9,158 $0.468$ $0.497$ $0.510$ $0.508$
ru 12,112 $0.551$ $0.580$ $0.563$ $0.630$
zh 7,085 $0.434$ $0.472$ $0.440$ $0.467$
ar 2,283 Low $0.438$ $0.465$ $0.45$ $0.445$
cy 1,241 $0.420$ $0.448$ $0.424$ $0.426$
et 1,782 $0.413$ $0.432$ $0.407$ $0.423$
id 1,243 $0.420$ $0.406$ $0.416$ $0.412$
ja 1,119 $0.417$ $0.412$ $0.411$ $0.422$
lv 2,337 $0.457$ $0.453$ $0.465$ $0.461$
mn 2,067 $0.429$ $0.441$ $0.419$ $0.437$
nl 7,108 $0.427$ $0.460$ $0.483$ 0.477
sl 1,843 $0.430$ $0.428$ $0.433$ $0.428$
sv 2,160 $0.454$ $0.448$ $0.472$ $0.454$
ta 1,358 $0.413$ $0.406$ $0.403$ $0.411$
tr 3,966 $0.480$ $0.519$ $0.484$ $0.514$
Table 7: Results of Table 1 in COMET score for each language pair. The best result in each group is denoted by ’_’.
Languages #Training data Resources Scratch ASR-PT Hu-Transformer DSU-Adapter
ca 95,854 High $18.24$ $25.66$ $24.24$ $25.98$
de 127,824 $13.97$ $21.06$ $19.28$ $21.32$
es 79,013 $22.37$ $29.8$ $26.64$ $29.34$
fr 207,372 $22.91$ $29.6$ $27.09$ $29.52$
fa 53,949 Mid $1.9$ $3.07$ $4.63$ $2.84$
it 31,698 $15.1$ $23.55$ $19.98$ $23.36$
pt 9,158 $7.01$ $11.2$ $13.13$ $11.69$
ru 12,112 $11.97$ $16.87$ $14.95$ $21$
zh 7,085 $3.58$ $6.39$ $4.4$ $5.67$
ar 2,283 Low $0.44$ $1.96$ $1.26$ $0.48$
cy 1,241 $0.64$ $1.24$ $1.21$ $0.72$
et 1,782 $0.49$ $1.46$ $0.61$ $0.93$
id 1,243 $0.31$ $0.67$ $1.27$ $0.79$
ja 1,119 $0.27$ $0.67$ $0.36$ $0.62$
lv 2,337 $1.22$ $1.65$ $2.63$ $2.06$
mn 2,067 $0.16$ $0.44$ $0.31$ $0.19$
nl 7,108 $2.14$ $6.95$ $10.75$ $8.46$
sl 1,843 $0.41$ $1.83$ $1.23$ $1.81$
sv 2,160 $0.9$ $1.28$ $3.4$ $1.76$
ta 1,358 $0.14$ $0.14$ $0.13$ $0.12$
tr 3,966 $1.67$ $3.5$ $3.04$ $3.14$
Table 8: Results of Table 1 in BLEU score for each language pair. The best result in each group is denoted by ’_’.

### v2

@@ 5.2.1 DSU-to-Trl: robust to tokenization?
When BPE is not applied on the DSU, those 6 DSU-to-Trl models have $6.48$$\pm$ $0.26$ BLEU. Despite having smaller model size (<$30$M), they are better than the Scratch model of $5.99$ BLEU.
When BPE is applied, the sequence length of DSU (DSU Length) could be shortened, which could in turn improve the performance, e.g. the best DSU-to-Trl model happens at configuration "$8$K" with $7.02$ BLEU. However, the DSU-to-Trl method is quite unstable to the use of BPE, as reflected by the $5.12$$\pm$ $0.83$ BLEU in the other 5 configurations. The correlation between the DSU sequence length, the source-target length ratio, and the ST performance is also not straightforward. For an example, the "$32$K" model (DSU length of $118$) is about $2$ BLEU behind to the "$8$K" model (DSU length of $150$). Therefore, applying BPE on the DSU for length reduction should remain cautious.
5.2.2 The DSU-Adapter is more robust
@@
Unlike DSU-to-Trl method, DSU-Adapter benefits more when BPE is applied to the DSU. Our proposed method has $8.86$$\pm$ $0.19$ BLEU (over the 6 corresponding configurations), as opposed to $8.17$$\pm$ $0.32$ BLEU when BPE is not applied. This observation is opposed to the DSU-to-Trl method which only scores $5.54$$\pm$ $1.07$ BLEU (with also larger variance) when BPE is applied on the DSU but $6.48$$\pm$ $0.26$ when BPE is not used. The improved mean score and its smaller variance suggests that the DSU-Adapter method is more (DSU) tokenization robust. We see this as a benefit of introducing the DSU, i.e., the SSS model knowledge, via PT rather than as model inputs.
On top of applying BPE on the DSU, using separate vocabularies in PT is preferred (the two yellow cells on Table 2) since it performs slightly better, and the DSU, which are not needed in the ST output, would not occupy the target vocabulary.
5.2.3 Ablation: initialisation in DSU-Adapter
@@
Having similar model sizes, e.g. about $50$M parameters (Table 2), DSU-Adapter is better than both EncDec-Init and Enc-Init methods. The translation performance in BLEU (averaged over the 12 vocabularies) is $8.51$$\pm$ $0.44$, $8.22$$\pm$ $0.51$, and $7.99$$\pm$ $0.61$ respectively. Encoder-initialization seems more crucial than decoder-initialization, as reflected by the fact that the best DSU-Adapter model comes from a combination with the weakest DSU-to-Trl model of $4.43$ BLEU.
5.3 Is CTC applicable also to DSU?
@@
Table 3: Effect of CTC on Fbk-to-DSU PT and/or ST FT to the DSU-Adapter method. All results are in BLEU and the best in each group is denoted by ’_’.
Similar to ST methods that use pretrained components, our method could be limited by the pretraining modality gap Liu et al. (2020); Le et al. (2023). Motivated by prior works, we investigate mitigating it with CTC. A crucial difference to the prior works is that our method uses DSU for pre-training rather than transcripts.
We thus study applying CTC in our method at different training stages. Owing to the large number of vocabulary configurations on Table 2, we only experiment with: 1) "No-BPE $1$K-$8$K", 2) "BPE $8$K", 3) "BPE $32$K" and 4) "BPE $32$K-$8$K". In each training stage, we report the effect of CTC to the ST performance (per resource group) by averaging the BLEU of these $4$ configurations.
Table 3 presents the analysis of applying CTC on our DSU-Adapter method. The training condition "Has CTC in DSU PT" refers to the case of applying CTC on the discrete speech units in Fbk-to-DSU pretraining, whereas "Has CTC in ST FT" refers to the case of applying CTC on the translations in ST finetuning, i.e., on the paired Fbk-Trl data. Our result shows that CTC helps on either stage, but the gain is $0.27$ BLEU more in ST finetuning. Using them jointly still helps, but the marginal gain is barely $0.03$ BLEU.
6 Limitations and future works
In the previous sections, we discuss the noticeable benefits of our DSU-pretraining method in creating a more compact ST model. In spite of this, there are several factors that are not thoroughly explored and could improve the model performance further:
@@ 6 Limitations and future works
We did not inspect the clustering size (fixed to $1,000$) and the number of training instances (only fixed to $98,000$) used in training the K-Means clustering model. Apart from tuning its hyper-parameters, using other techniques, such as residual vector quantisation Zeghidour et al. (2021); Défossez et al. (2022) and multiple codebooks Guo et al. (2023), might bring better improvement.
@@
A stronger pretrained decoder
Apart from strengthening the encoder, the DSU-to-Trl model and hence its decoder (used in initialisation) could also be improved, e.g. via back-translation, up-sampling the textual sequence Yan et al. (2024) and pretraining with more text data, while maintaining the small decoder size.
Further analyses
In addition to improving our pretraining method for better model compactness, there are other related research directions worth further analyzing. One direction would be how, in terms of acoustic pretraining, DSU compared with transcripts (if available in that language) over different data scales. Another interesting research direction would be the comparison and analysis of using DSU or dense features in a large pretrained model setting, such as Whisper Radford et al. (2023) and Large Language Models.
7 Conclusion
In this paper, we consider a memory-constrained setting for ST. Our proposed method uses DSU in the form of pretraining to distil the knowledge from the Self-Supervised Speech model to creating more compact Speech-to-text Translation. Our compact model, i.e., the DSU-Adapter, shows strong and consistent improvements in three evaluation metrics over the baselines. In contrast to using DSU as model inputs, our method does not require quantization modules in inference and shows stronger robustness to the DSU tokenization. Finally, our method requires no transcripts, making it also suitable for low-resource setting.
Figure 3: Results in BLEU, chrF and COMET-22-DA on each language pair of CoVoST-2 (X-En).